OpenStudy (anonymous):

I really need some help on this question. It wants me to write this equation, 3x^2-9x+38=2x^2+18, in standard form (ax^2+bx+c=0, I think) and then factor the resulting equation.

OpenStudy (anonymous):

I got it

OpenStudy (anonymous):

3x^2-2x^2-9x=18-38 x^2-9x=-20

OpenStudy (anonymous):

@hsedward43